find the basis and the dimension of the solution space W of the following system

X1 + 2X2 - 2X3 + 2X4 - X5 = 0

X1 + 2X2 - X3 + 3X4 - 2X5 = 0

2X1 + 4X2 - 7X3 +X4 - X5 = 0


1
Expert's answer
2021-08-23T17:04:58-0400

Let us find the basis and the dimension of the solution space "W" of the following system


"\begin{cases}
X_1 + 2X_2 - 2X_3 + 2X_4 - X_5 = 0\\

X_1 + 2X_2 - X_3 + 3X_4 - 2X_5 = 0\\

2X_1 + 4X_2 - 7X_3 +X_4 - X_5 = 0
\end{cases}"


After substracting from the second equation the first equation, and after multiplying the first equation by "-2" and adding to the third equation, we get the following equivalent system:


"\begin{cases}
X_1 + 2X_2 - 2X_3 + 2X_4 - X_5 = 0\\

X_3 + X_4 - X_5 = 0\\

-3X_3 -3X_4 +X_5 = 0
\end{cases}"


Let us multiply the second equation by 3 and add to the third equation:


"\begin{cases}
X_1 + 2X_2 - 2X_3 + 2X_4 - X_5 = 0\\

X_3 + X_4 - X_5 = 0\\

-2X_5 = 0
\end{cases}"


It follows that


"\begin{cases}
X_1 =- 2X_2 +2X_3 -2X_4 \\

X_3 =- X_4\\

X_5 = 0
\end{cases}"


and hence


"\begin{cases}
X_1 =- 2X_2 -4X_4 \\

X_3 =- X_4\\

X_5 = 0
\end{cases}"


We conclude that


"W=\{(X_1,X_2,X_3,X_4,X_5)\in\R^5: X_1 =- 2X_2 -4X_4,\ X_3 =- X_4, \ X_5 = 0\}\\
=\{(- 2X_2 -4X_4,X_2,-X_4,X_4,0)\in\R^5: X_2, X_4 \in\R\}\\
=\{X_2(- 2,1,0,0,0)+X_4(-4,0,-1,1,0)\in\R^5: X_2, X_4 \in\R\}."


Therefore, "\{(- 2,1,0,0,0),(-4,0,-1,1,0)\}" is a basis for the solution space "W" of the system, and hence "dim\ W=2."